OpenStudy (anonymous):

when the following expressions are for example 3to the zero power does it equal =0

OpenStudy (australopithecus):

x^(0) = 1 anything to the exponent zero is equal to 1

OpenStudy (australopithecus):

Note that x^(-1) = 1/x and x^(1) = x so x/x = x^(1)x^(-1) = x^(0) = 1 this makes sense because x/x = 1 remember add exponents when multiplying

OpenStudy (australopithecus):

thus 1 + (-1) = 0

OpenStudy (australopithecus):

that is why any number to the power of zero is equal to 1

OpenStudy (anonymous):

so 3^(0) =1

OpenStudy (australopithecus):

3^(-1) = 1/3 3^(1) = 3 thus 3/3 = 3^(1)3^(-1) = 3^(0) = 1 3/3 = 1

OpenStudy (australopithecus):

thus 3^(0) =1 anything to the power of zero equals 1 (x-1)^(0) = 1 34343^(0) = 1 etc..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Not trying to be nitpicky, but 0^0 is indeterminate. Everything else is 1.

OpenStudy (australopithecus):

not nitpicking bmp of course there are exceptions to the rule (1/0)^(0) =/= 1

OpenStudy (australopithecus):

but that is a more relevant case and I'm glad you brought it up

OpenStudy (anonymous):

? I just started learning this yesterday so in all due respect im confused about indeterminate etc

OpenStudy (australopithecus):

dont worry about it deenarae just remember what I showed you and remember that 0^(0) is not solvable/cannot be determined.

OpenStudy (anonymous):

i thought you said it equals 1

OpenStudy (anonymous):

@Australopithecus Thanks for not taking it badly. I wasn't trying to be a smart retrice just trying to be more complete. :-)

OpenStudy (australopithecus):

everything equals 1 but you cannot divide by 0 thus if you set something up like this 0/0 =/= 1 because you are not allowed to divide by zero in mathematics.

OpenStudy (australopithecus):

the reason you cannot divide by zero is because the concept of nothing going into something doesn't make sense

OpenStudy (anonymous):

so its undefined correct

OpenStudy (australopithecus):

correct

OpenStudy (anonymous):

ok thanks much makes sense

OpenStudy (australopithecus):

everything else is equal to 1

OpenStudy (anonymous):

so the zero power of 3 =1

OpenStudy (australopithecus):

yes

OpenStudy (anonymous):

k

OpenStudy (australopithecus):

if you have any other questions please feel free to ask

OpenStudy (anonymous):

ty